OpenStudy (anonymous):

Divide: (3x^3+2x^2-x-10) / (x+2)

OpenStudy (anonymous):

Quotient is \[3x ^{2} - 4x + 7\] and remainder is -24

OpenStudy (anonymous):

How did you get that though?

OpenStudy (anonymous):

Is it correct???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay I will show the working It is by the LONG DIVISION METHOD as follows: 3x^2 - 4x + 7 ------------------------ x+2 | 3x^3 + 2x^2 - x - 10 -3x^3 - 6x^2 ---------------------- -4x^2 - x - 10 +4x^2 +8x --------------- 7x - 10 -7x - 14 ---------- - 24 Is that clear enough???

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Got a medal for me??

OpenStudy (anonymous):

Mathematica confirms that Harkirat is correct. \[\text{PolynomialQuotient}\left[-10-x+2 x^2+3 x^3,2+x,x\right]=7-4 x+3 x^2 \] \[\text{PolynomialRemainder}\left[-10-x+2 x^2+3 x^3,2+x,x\right] =-24\] Congratulations Harkirat. It has been alot years since I've seen that done by hand. Very interesting.